I am unable to solve
m=6 over7,(4,-9)

Answer by mananth(16946)

m= 6/ 7

Plug value of the slope and point (4, -9) in
Y = m x + b
-9.00 = 3 3/7 + b
b= -9 - 24/7
b= -87/ 7
So the equation will be
Y = 6/7 x -87/7
m.ananth@hotmail.ca